# When Debate Fails: Bias Reinforcement in Large Language Models

### v1

@@ 1 Introduction
  • We identify the bias reinforcement problem in Multi-Agent Debate, showing that MAD strengthens both correct and incorrect reasoning rather than inherently improving it.
@@ 3.1 Overview
We introduce MetaNIM Arena, illustrated in Figure 8. It features four impartial combinatorial games, meaning both players share identical moves at each state, all information is fully observable, and each game terminates in a finite number of moves. By merging combinatorial game theory with adversarial play, MetaNIM Arena serves as a benchmark for debate-based strategic reasoning in LLMs.
Dataset and Simulator. Key game situations are systematically collected into a dataset, each accompanied by an optimal action, enabling LLMs to be tested for decision-making accuracy. These results appear in Table 3, with details in Appendix A.7. Separately, MetaNIM Arena can function as a simulator: the model encounters an adaptive opponent, so each trajectory depends on both the agent’s and the opponent’s actions. Here, binary win/loss outcomes allows evaluation by win rate. We demonstrate this approach in Table 4, with further explanation in Appendix A.8. We refer the Appendix B for details of opponent modelling.
@@ 3.2 Why MetaNIM Arena?
2. Clear Optimality Criterion: By the Sprague-Grundy Theorem, these games admit an optimal strategy. MetaNIM Arena thus measures how closely a model’s reasoning aligns with that strategy, instead of relying on approximate metrics like BLEU or perplexity.
We also assert that MetaNIM Arena naturally supports Reinforcement Learning framework. By providing a binary win/loss signal and structuring gameplay as a Markov Decision Process (MDP), it enables iterative strategy refinement—an advantage often absent in static benchmarks.
@@ 3.3 Game Variants
Here, we introduce four settings of MetaNIM ArenaDetailed explanation and theoretically determined winning strategies for these games are provided in Appendix A.
@@ 3.4 Combinatorial Games: Theory and Strategy
$G(S)\;=\;\mathrm{mex}\{\,G(S^{\prime})\;|\;S^{\prime}\!\text{ is reachable from }S\,\}\,,$
where $\mathrm{mex}(X)$ is the smallest nonnegative integer not in $X$.
Note that the Grundy number is well-defined for every impartial combinatorial game, since the game’s state space forms a DAG. In many combinatorial games, direct enumeration of all possible move sequences is computationally infeasible. However, with Sprague-Grundy Theorem, we can easily calculate Grundy Numbers on complex games. See Appendix A.1 for further discussions.
Optimal Strategy. When $G(S)\neq 0$, there is at least one move to a successor $S^{\prime}$ with $G(S^{\prime})=0$, forcing the opponent into a losing position. Conversely, if $G(S)=0$, all successor states have $G(S^{\prime})\neq 0$. Because the game DAG is finite and acyclic, repeatedly applying “move to $G=0$” (or avoiding it) ensures a forced result under optimal play. See Appendix A.2 for more details, including the misère variant where taking the last object loses.
@@ 4.1 Bias Reinforcement in MAD
(a) GPT-4o Wrong Bias Good Bias
(20, 19) (12, 4) (7, 4) (15, 10) (16, 8) (7, 7)
Standard 0.700 0.675 0.600 0.525 0.725 0.850
+ After MAD 0.900 0.750 0.700 0.750 0.750 0.900
(b) GPT-4o-mini Wrong Bias Good Bias
(18, 4) (12, 6) (10, 4) (15, 10) (7, 2) (15, 2)
Standard 0.700 0.875 0.600 0.975 0.950 0.975
+ After MAD 0.850 0.950 0.750 1.000 0.950 1.000
(c) GEMINI-1.5-pro Wrong Bias Good Bias
(20, 19) (12, 4) (4, 4) (15, 10) (10, 4) (16, 8)
Standard 0.650 0.600 0.600 0.800 0.625 0.675
+ After MAD 0.700 0.650 0.800 0.800 0.800 0.700
(d) GEMINI-1.5-flash Wrong Bias Good Bias
(12, 6) (12, 4) (7, 4) (15, 4) (20, 19) (7, 7)
Standard 0.800 0.700 0.525 0.750 0.500 0.750
+ After MAD 0.850 1.000 0.750 0.750 0.650 1.000
In NIM: (Figure 1) We observed that instead of refining reasoning towards optimal strategies, MAD reinforces a model’s inherent biases, leading the debate process to align with strongly consistent incorrect responses. To systematically analyze this effect, we identified states where each model exhibits strong consistency. We then conducted multi-agent debates using two identical agents initialized with the same model.
@@
In our setup, we collected 40 trials per state, with each agent generating 20 responses at a fixed sampling temperature of 0.7. The same temperature was maintained during the MAD process to ensure consistent evaluation. Initial action distributions are shown in light red, while blue bars represent distributions after three rounds of debate. Ideally, if MAD functioned as an effective self-correction mechanism, the distribution would shift toward optimal actions.
@@
However, we found the opposite: regardless of whether the initial reasoning was correct, the debate process consistently amplifies pre-existing biases rather than mitigating them. For instance, in the top-left graph of Figure 1, the GPT-4o-mini model initially selects a strongly consistent but suboptimal action (Action 3) 82.5% of the time, while only 15.0% of responses align with the optimal strategy. After multiple rounds of debate, instead of converging towards the correct decision, the model further entrenches its bias, increasing the frequency of the incorrect action to 90.0%. Notably, the proportion of optimal responses decreased even further, reinforcing the idea that MAD serves as an amplifier of existing biases rather than a corrective mechanism.
This phenomenon is not unique to GPT models but is also evident in the GEMINI series (bottom row of Figure 1). Despite architectural differences, GEMINI exhibited the same tendency. Furthermore, we observed that this reinforcement occurs regardless of whether the bias is “wrong” (leading to suboptimal moves) or “good” (aligning with optimal play). As illustrated, even in states where strong consistency initially aligns with optimal actions (good bias cases), the debate process did not introduce additional strategic insights but simply reinforces what is already dominant. In Figure 2, we illustrate in detail how two LLMs, after receiving the current state, produce initial responses and then proceed through a debate process. Notably, they converged on a strongly consistent line of reasoning, which may not necessarily be correct. We observed that only after the first round of debate do the two LLMs reach a consensus.
In Fibonacci: (Table 2) Following our analysis in the NIM game, we extended our investigation to the Fibonacci game, a more complex setting where move constraints and dynamic interactions shape strategic decision-making. We identified states where models exhibit strong consistency and examined whether debate reinforces or mitigates initial biases. As in NIM, we categorize states into those where strong consistency aligns with optimal strategies and those where it leads to suboptimal choices. The MAD framework is then applied to analyze how action distributions evolve post-debate (Table 2 (a), (b) for GPT models and (c), (d) for GEMINI models).
Our results confirm a consistent bias reinforcement effect: rather than converging toward optimal strategies, MAD causes models to reinforce their strongly consistent responses, regardless of correctness. As shown in Table 2, the proportion of responses adhering to the initially strongly consistent action increases after debate, indicating further entrenchment of bias. Our analysis shows that after the debate process, response reinforcement increases by 9.17% in GPT models and 12.29% in GEMINI models, indicating a notable shift towards initially dominant responses across both model families.
Figure 2: An example demonstrating how the debate process converges to a biased outcome. We observed that bias reinforcement occurs in the first debate. Blue text indicates the correct reasoning and orange text indicates the strong consistent (biased) reasoning. We skipped the second debate process because it follows a process nearly identical to those of the first and third debates. We utilize GPT-4o-mini as the debate agent.
@@
Figure 3: Decline in optimal actions over debate rounds, demonstrating the convergence toward consistently biased reasoning. This observation underscores the difficulty of addressing inherent biases in LLMs through MAD.
@@ 4.2 Optimal Inputs Do Not Reduce Bias in MAD
To analyze the extent of bias reinforcement in MAD, we conducted an additional experiment using GPT-4o-mini in the NIM game (remaining items: 5), as shown in Figure 3. We first collected 20 responses demonstrating strong consistency, where the model repeatedly selects the same biased action (Action 2) with high probability. Additionally, we curated another set of 20 responses exhibiting optimal strategic reasoning, where 80% of actions align with the game-theoretically optimal move (Action 1). We then introduced these responses into a multi-agent debate setting using GPT-4o-mini, pairing each strongly consistent response with an optimal response, and observed how the model’s reasoning evolves over multiple rounds of debate (denoted the game situation in detail in Appendix  E).
Our findings reveal that rather than converging toward more optimal decision-making, the debate systematically reinforces the model’s pre-existing biases. Initially, the curated dataset contained 80% optimal responses, yet after a single round of debate, the model predominantly aligned with the biased responses, reducing the proportion of optimal decisions. As the debate progressed, this effect intensified, with strong consistency actions increasingly dominating. That is, even when explicitly injecting high-quality optimized responses, the debate process does not steer the model toward better reasoning. Instead, it systematically aligns with its pre-existing consistency patterns, reinforcing suboptimal but frequent choices.
This limitation in MAD highlights its tendency to amplify biases rather than correct them when strong consistency is present. However, applying DReaMAD with curated optimal responses significantly improved the preservation of correct reasoning, mitigating bias reinforcement. This underscores the need for mechanisms that introduce diverse perspectives beyond internal debate, which we further discuss in §5.
@@ 4.3 Lack of Reasoning Diversity in MAD
While MAD is designed to refine reasoning by allowing multiple agents to engage in discussion, its effectiveness is fundamentally limited by a lack of true diversity (§2.3). To address this limitation, Chen et al. (2023b) propose a multi-model debate framework that leverages multiple different models to generate diverse reasoning paths. However, our research focuses on self-correction methods that utilize a single model. We observe that even within the same model, slight variations in the prompts can lead to significantly different outputs (demonstrated in Appendix D). For instance, simply including or omitting the word Fibonacci in the prompt results in distinct reasoning trajectories. This suggests that diversity in reasoning can be enhanced by strategically modifying prompts within the same model, providing a practical alternative to multi-model debate frameworks.
5 DReaMAD: Diverse Reasoning via Multi-agent Debate
Figure 4: DReaMAD framework in MAD method. DReaMAD improves LLM reasoning by combining Strategic Prior Knowledge Elicitation and Perspective Diversification. In the first stage, the model reinterprets the problem and formulates high-level strategies to reduce bias. In the second stage, multiple agents adopt distinct viewpoints, engage in structured debate, and refine their conclusions to enhance decision-making.
@@ 5 DReaMAD: Diverse Reasoning via Multi-agent Debate
5.1 Strategic Prior Knowledge Elicitation
5.2 Perspective Diversification
Building on MAD, DReaMAD mitigates argument homogenization by ensuring each agent adopts a distinct viewpoint prior to the debate. This approach is inspired by the Learning from Multiple Approaches concept in education theory (Council et al., 2005; Cleaves, 2008), which suggests that individuals improve their problem-solving skills by exploring multiple representations of the same problem. Analogously, this module ensures that each agent is given differentiated initial prompts that guide reasoning along distinct strategic trajectories. By self-customizing initial prompts for each model instance, DReaMAD encourages unique strategic perspectives and reduces the risk of bias reinforcement (Figure 4-2) as demonstrated in Figure 5 and Figure 6. In Figure 5, although each agent receives the same initial prompt, they independently generate different optimal prompts, leading to distinct distributions of reasoning. This process reduces each agent’s bias and fosters a more robust debate. Remarkably, even in a state with very strong consistency, the discussion converges well toward the correct reasoning direction. Here, we set temperature to be 0.7 for diversity.
6 Experiments
We utilized the new benchmark MetaNIM Arena as both our dataset and simulator. Our investigation focuses on three key questions: (1) Does our approach improve reasoning quality compared to existing prompting techniques? (2) Does our approach prove its strategic reasoning quality under adversarial decision-making environments? (3) Does generating diverse prompts contribute to better decision-making within the debate framework?
To address these questions, we compared DReaMAD against standard prompting methods, ReAct (Yao et al., 2023), Chain-of-Thought (CoT), Self-Consistency (Wang et al., 2022), Self-Refinement (Madaan et al., 2024), and Multi-Agent Debate (MAD; Du et al. (2023) (standard and CoT prompts is explained in Appendix B). Our algorithm is built upon the work of Du et al. (2023), extending MAD to incorporate structured self-prompt refinement and perspective diversification. For self-refinement, we follow the methodology of Madaan et al. (2024), applying three iterative refinement steps. Similarly, for MAD, we conducted up to three rounds of debate, following Du et al. (2023), with the process terminating early if a consensus is reached before the final round. Additionally, we analyzed whether different LLM architectures, including GPT and GEMINI models, exhibit similar performance improvement through DReaMAD.
6.1 Does DReaMAD Improve Reasoning Quality?
Figure 6: Effect of Perspective Diversification with two different prompts. (+)MAD vs DReaMAD. We run LLMs total 50 runs for each simulator and report average win-rate.

### v2

# Understanding Bias Reinforcement in LLM Agents Debate

@@ 1 Introduction
  • We identify the bias reinforcement in Multi-Agent Debate, showing that MAD strengthens both correct and incorrect reasoning rather than inherently improving it.
@@ 2.2 Prompt Engineering and Self-Correction in LLMs
Figure 1: An example demonstrating how the debate process converges to a biased outcome. We observed that bias reinforcement occurs in the first debate. Blue text indicates the correct reasoning and orange text indicates the strong consistent (biased) reasoning. The second debate is omitted, as its procedure replicates the first and third; all debates use GPT-4o-mini as the debating agent.
@@ 3.1 Overview
We introduce MetaNIM Arena, illustrated in Figure 7. It features six impartial combinatorial games, meaning both players share identical moves at each state, all information is fully observable, and each game terminates in a finite number of moves. By merging combinatorial game theory with adversarial play, MetaNIM Arena serves as a benchmark for debate-based strategic reasoning in LLMs.
Dataset and Simulator. Key game situations are systematically collected into a dataset, each accompanied by an optimal action, enabling LLMs to be tested for decision-making accuracy. These results appear in Table 3, with details in Appendix A.3. Separately, MetaNIM Arena can function as a simulator: the model encounters an adaptive opponent, so each trajectory depends on both the agent’s and the opponent’s actions. Here, binary win/loss outcomes allows evaluation by win rate. We demonstrate this approach in Table 4, with further explanation in Appendix A.4. We refer the Appendix B for details of opponent modeling.
GPT-4o Wrong Bias Good Bias
(20, 19) (12, 4) (7, 4) (15, 10) (16, 8) (7, 7)
Standard 0.700 0.675 0.600 0.525 0.725 0.850
+ After MAD 0.900 0.750 0.700 0.750 0.750 0.900
GPT-4o-mini Wrong Bias Good Bias
(18, 4) (12, 6) (10, 4) (15, 10) (7, 2) (15, 2)
Standard 0.700 0.875 0.600 0.975 0.950 0.975
+ After MAD 0.850 0.950 0.750 1.000 0.950 1.000
GEMINI-1.5-pro Wrong Bias Good Bias
(20, 19) (12, 4) (4, 4) (15, 10) (10, 4) (16, 8)
Standard 0.650 0.600 0.600 0.800 0.625 0.675
+ After MAD 0.700 0.650 0.800 0.800 0.800 0.700
GEMINI-1.5-flash Wrong Bias Good Bias
(12, 6) (12, 4) (7, 4) (15, 4) (20, 19) (7, 7)
Standard 0.800 0.700 0.525 0.750 0.500 0.750
+ After MAD 0.850 1.000 0.750 0.750 0.650 1.000
@@ 3.2 Why MetaNIM Arena?
2. Clear Optimality Criterion: By the Sprague-Grundy Theorem, these games admit an optimal strategy. MetaNIM Arena thus measures how closely a model’s reasoning aligns with that strategy, instead of relying on approximate metrics like BLEU or perplexity. We also assert that MetaNIM Arena naturally supports reinforcement learning framework. By providing a binary win/loss signal and structuring gameplay as a Markov Decision Process (MDP), it enables iterative strategy refinement—an advantage often absent in static benchmarks.
@@ 3.3 Game Variants
Here, we introduce four settings of MetaNIM Arena. Detailed explanation and theoretically determined winning strategies for these games are provided in Appendix A.
@@
  • Corner Queen: On a rectangular board, two players take turns moving a queen left, down, or diagonally down-left. The first to reach the bottom-left corner wins. Strategy lies in limiting the opponent’s options while advancing.
@@ 3.4 Combinatorial Games: Theory and Strategy
$G(S)\;=\;\mathrm{mex}\{\,G(S^{\prime})\;|\;S^{\prime}\!\text{ is reachable from }S\,\}\,.$
Here, $\mathrm{mex}(X)$ is the smallest nonnegative integer not in $X$.
Note that the Grundy number is well-defined for every impartial combinatorial game, since the game’s state space forms a DAG. In many combinatorial games, direct enumeration of all possible move sequences is computationally infeasible. However, with Sprague-Grundy Theorem, we can easily calculate Grundy Numbers on complex games. See Appendix A.1.1 for further discussions.
Optimal Strategy. When $G(S)\neq 0$, there is at least one move to a successor $S^{\prime}$ with $G(S^{\prime})=0$, forcing the opponent into a losing position. Conversely, if $G(S)=0$, all successor states have $G(S^{\prime})\neq 0$. Because the game DAG is finite and acyclic, repeatedly applying “move to $G=0$” (or avoiding it) ensures a forced result under optimal play. See Appendix A.1.2 for more details, including the misère variant where taking the last object loses.
@@ 4.1 Bias Reinforcement in MAD
In NIM: (Figure 2) We find that MAD amplifies models’ pre-existing biases rather than refining their reasoning. To investigate this, we first identify game states where each model exhibits strong consistency, i.e., consistently selecting the same action across multiple trials. For each such state, we conduct multi-agent debates using two identical agents instantiated from the same model.
@@
Each agent generates 20 responses (40 per state) at a fixed temperature of 0.7, maintained throughout the debate. Initial action distributions (light red) are compared against post-debate distributions after three rounds (blue). If MAD functioned as a self-correction mechanism, we would expect the distribution to shift toward the optimal action.
@@
However, we found the opposite: regardless of whether the initial reasoning was correct, the debate process consistently amplifies pre-existing biases rather than mitigating them. For instance, in Figure 2 (top-left), GPT-4o-mini initially selects a suboptimal action (Action 3) 82.5% of the time. After the debate, this frequency increases to 90.0%, while the proportion of optimal responses drops further. Rather than correcting errors, the debate reinforces strongly consistent—but incorrect—responses. This pattern persists across model families. The GEMINI models (bottom row of Figure 2) exhibit similar behavior, regardless of whether the initial bias aligns with optimal play (“good bias”) or not (“wrong bias”). In both cases, MAD strengthens the dominant trajectory without introducing new strategic insight.
Figure 1 provides a detailed view: two LLM agents receive the same input and begin debating. Despite initial divergence, they converge quickly—after the first round—on a shared line of reasoning. Crucially, this convergence occurs even when the initial consensus is incorrect, illustrating that MAD often serves as an amplifier of bias rather than a correction mechanism.
In Fibonacci: (Table 2) Extending our NIM analysis, we evaluate MAD’s effect in the Fibonacci game, a more complex setting with move constraints and dynamic interactions. As in NIM, We identify states exhibiting strong consistency and categorize them into two groups: those where consistent responses align with the optimal strategy, and those where they do not. We then apply MAD to examine how response distributions evolve post-debate.
Consistent with the NIM results, MAD reinforces the model’s initial biases in Fibonacci as well. As shown in Table 2, the frequency of initially consistent actions increases after debate, regardless of whether those actions are optimal. On average, reinforcement rises by 9.17% in GPT models and 12.29% in GEMINI models, indicating a systematic amplification of dominant reasoning patterns across architectures.
@@
Figure 3: Decline in optimal actions over debate rounds, demonstrating the convergence toward consistently biased reasoning.
@@ 4.2 Optimal Inputs Do Not Reduce Bias in MAD
Figure 4: DReaMAD framework. DReaMAD improves LLM reasoning by combining Strategic Prior Knowledge Elicitation and Perspective Diversification. In the first stage, the model reinterprets the problem and formulates high-level strategies to reduce bias. In the second stage, multiple agents adopt distinct viewpoints, engage in structured debate, and refine their conclusions to enhance decision-making.
To further assess the bias reinforcement in MAD, we conduct an experiment on the NIM game (5 items remaining) using GPT-4o-mini, illustrated in Figure 3. We prepare two sets of 20 responses: one exhibiting strong consistency toward a biased action (Action 2), and another consisting of optimal responses, where 80% of actions correspond to the game-theoretic best move (Action 1). We then introduced these responses into a multi-agent debate setting using GPT-4o-mini, pairing each strongly consistent response with an optimal response, and observed how the model’s reasoning evolves over multiple rounds of debate (denoted the game situation in detail in Appendix  E).
Surprisingly, the debate fails to leverage the high-quality input. Initially, the curated dataset contained 80% optimal responses, yet after a single round of debate, the model predominantly aligned with the biased responses. As debate rounds proceed, this effect intensifies: the influence of the optimal input diminishes, while the initially frequent—but suboptimal—choice becomes dominant. These results demonstrate that MAD is not only ineffective at correcting bias, but can systematically aligns with its pre-existing consistency patterns, reinforcing suboptimal but frequent choices.
In contrast, applying DReaMAD with the curated optimal responses preserves correct reasoning and mitigating bias reinforcement. This underscores the need for mechanisms that introduce diverse perspectives beyond internal debate, which we further discuss in §5.
@@ 4.3 Lack of Reasoning Diversity in MAD
While MAD is designed to refine reasoning through agent interaction, its effectiveness is fundamentally constrained by a lack of genuine diversity (§2.3). To address this limitation, Chen et al. (2023b) propose a multi-model debate framework that combines outputs from different models to encourage diverse reasoning. In contrast, our approach focuses on single-model self-correction.
Interestingly, we observe that even within a single model, small variations in prompts can induce markedly different reasoning paths (Appendix D). For example, including or omitting the word Fibonacci leads to distinct strategies in the same task. This suggests that diversity in reasoning can be enhanced by strategically modifying prompts within the same model, providing a practical alternative to multi-model debate frameworks.
5 DReaMAD: Diverse Reasoning via MAD
@@
5.1 Strategic Prior Knowledge Elicitation
5.2 Perspective Diversification
Building on MAD, DReaMAD mitigates argument homogenization by ensuring each agent adopts a distinct viewpoint prior to the debate. This approach is inspired by the Learning from Multiple Approaches concept in education theory (Council et al., 2005; Cleaves, 2008), which suggests that individuals improve their problem-solving skills by exploring multiple representations of the same problem. Analogously, this module ensures that each agent is given differentiated initial prompts that guide reasoning along distinct strategic trajectories. By self-customizing initial prompts for each model instance, DReaMAD encourages unique strategic perspectives and reduces the risk of bias reinforcement (Figure 4-2) as demonstrated in Figure 5 and Figure 6 (left). In Figure 5, although each agent receives the same initial prompt, they independently generate different optimal prompts, leading to distinct distributions of reasoning. This process reduces each agent’s bias and fosters a more robust debate. Remarkably, even in a state with very strong consistency, the discussion converges well toward the correct reasoning direction. Here, we set temperature to be 0.7 for diversity.
6 Experiments
LLM Models Prompting Methods NIM Fibonacci Kayles Chomp C.Queen
Normal Misère Normal Misère Single 2 Rows Rectangular Square Normal
GEMINI-1.5-flash Standard Prompting 0.32 0.54 0.16 0.10 0.54 0.56 0.78 0.18 0.46
+ ReAct 0.10 0.68 0.16 0.76 0.50 0.30 0.42 0.12 0.46
+ Self-Refinement 0.14 0.66 0.18 0.36 0.50 0.46 0.46 0.16 0.42
+ Self-Consistency 0.04 0.28 0.28 0.86 0.30 0.24 0.74 0.0 0.34
+ MAD 0.06 0.30 0.12 0.78 0.54 0.20 0.74 0.14 0.58
+ MAD2 0.26 0.26 0.08 0.06 0.44 0.36 0.68 0.10 0.58
+ DReaMAD 0.38 0.84 0.16 0.94 0.58 0.62 0.60 0.22 0.74
GPT-4o-mini Standard Prompting 0.38 0.54 0.22 0.28 0.46 0.48 0.46 0.38 0.12
+ ReAct 0.22 0.68 0.20 0.34 0.40 0.26 0.58 0.34 0.28
+ Self-Refinement 0.22 0.70 0.18 0.50 0.46 0.52 0.52 0.44 0.24
+ Self-Consistency 0.14 0.52 0.34 0.46 0.32 0.20 0.54 0.26 0.30
+ MAD 0.28 0.62 0.22 0.64 0.42 0.28 0.52 0.56 0.44
+ MAD2 0.34 0.20 0.18 0.18 0.50 0.34 0.44 0.90 0.14
+ DReaMAD 0.98 0.74 0.54 0.72 0.68 0.84 0.64 0.22 0.76
We utilized the benchmark MetaNIM Arena as both our dataset and simulator, as it provides a controlled environment for evaluating reasoning under grounded strategic tasks. Our investigation focuses on three key questions: (1) Does our approach improve reasoning quality compared to existing prompting techniques? (2) Does our approach prove its strategic reasoning quality under adversarial decision-making environments? (3) Does generating diverse prompts contribute to better decision-making within the debate framework?
To address these questions, we compare DReaMAD with standard prompting methods including ReAct (Yao et al., 2023), Chain-of-Thought (CoT), Self-Consistency (Wang et al., 2022), Self-Refinement (Madaan et al., 2024), and Multi-Agent Debate (MAD (Du et al., 2023), MAD2 (Liang et al., 2023)). The details of standard and CoT prompts are provided in Appendix B.
Our method builds on the MAD framework by Du et al. (2023), augmenting it with structured self-prompt refinement and perspective diversification. For self-refinement, we follow the methodology of Madaan et al. (2024), applying three iterative refinement steps. Similarly, for MAD, we conducted up to three rounds of debate, following Du et al. (2023), with the process terminating early if a consensus is reached before the final round.
We also investigate whether the observed improvements generalize across different LLM architectures, including both GPT and GEMINI models.
Table 5: Accuracy (%) on math-reasoning benchmarks. Columns group the five algorithms for each backbone model. Bold indicates the best algorithm for a given (dataset, model) pair.
Dataset GPT-o3-mini GPT-4o
ReAct Self-Refinement Self-Consistency MAD DReaMAD ReAct Self-Refinement. Self-Consistency. MAD DReaMAD
AIME 2024 76.7 73.3 86.7 73.3 90.0 0.0 3.3 10.0 3.3 10.0
AMC 2023 97.5 100 100 100 100 60.0 52.5 52.5 60.0 62.5
Table 6: Accuracy (%) on CommonsenseQA dataset. Bold indicates the best performance. We abbreviate Self-Refinement as Self-Refine. and Self-Consistency as Self-Consist.
Model ReAct Self-Refine. Self-Consist. MAD DReaMAD
GPT-4o 83.6 49.2 83.6 82.0 84.4
GPT-4o-mini 78.7 61.5 78.7 75.4 79.5
6.1 Does DReaMAD Improve Reasoning Quality?
Figure 6: Effect of perspective diversification. Left: Average win rate of (+)MAD and DReaMAD on NIM and Fibonacci (Normal and Misère variants), aggregated over 50 simulations per setting. Right: Accuracy on the Fibonacci benchmark with GPT-4o across different sampling temperatures (15 runs each, 95% CI). Higher temperatures yield greater prompt diversity, leading to improved accuracy.